Suppose that a pair of fair dice is thrown 1 time. Let the random variable X be the absolute difference of the dots that appear.


1) - What are the possible values of X?

2) - What is the sample space?

3) - Find the probability functions of X.

4) - Find the probability mass function

5) - Draw the graph for the probability mass function




1
Expert's answer
2022-05-24T17:54:21-0400
"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c:c:c}
 & 1 & 2 & 3 & 4 & 5 & 6 \\ \hline
 1 & 0 & 1 & 2 & 3 & 4 & 5 \\
 \hdashline
 2 & 1 & 0 & 1 & 2 & 3 & 4 \\
 \hdashline
 3 & 2 & 1 & 0 & 1 & 2 & 3 \\
 \hdashline
 4 & 3 & 2 & 1 & 0 & 1 & 2 \\
 \hdashline
 5 & 4 & 3 & 2 & 1 & 0 & 1 \\
 \hdashline
 6 & 5 & 4 & 3 & 2 & 1 & 0 \\
 \hdashline
\end{array}"


1) The possible values of "X" are "0,1, 2, 3, 4," and "5."


2)

"S=\{0,1,2,3,4,5\}"

3)


"\def\arraystretch{1.5}
 \begin{array}{c:c}
 x & 0 & 1 & 2 & 3 & 4 & 5 \\ \hline
 p(x) & 6/36 & 10/36 & 8/36 & 6/36 & 4/36 & 2/36 \\

\end{array}"


4)


"f(x) = \begin{cases}
 6/36 &\ x=0 \\
 10/36&\ x=1 \\
 8/36&\ x=2 \\
 6/36&\ x=3 \\
 4/36&\ x=4 \\
 2/36&\ x=5 \\
\end{cases}"

5)